# How Not to Detect Prompt Injections with an LLM

### v2

@@ Abstract
LLM-integrated applications and agents are vulnerable to prompt injection attacks, where adversaries embed malicious instructions within seemingly benign input data to manipulate the LLM’s intended behavior. Recent defenses based on known-answer detection (KAD) scheme have reported near-perfect performance by observing an LLM’s output to classify input data as clean or contaminated. KAD attempts to repurpose the very susceptibility to prompt injection as a defensive mechanism. In this work, we formally characterize the KAD scheme and uncover a structural vulnerability that invalidates its core security premise. To exploit this fundamental vulnerability, we methodically design an adaptive attack, DataFlip. It consistently evades KAD defenses, achieving detection rates as low as $1.5\%$ while reliably inducing malicious behavior with success rates of up to $88\%$—all without requiring white-box access to the LLM or any optimization procedures. We release our evaluation code at [1].
1 Introduction
@@
Large Language Models (LLMs) enable modern agentic systems [2] and AI-driven applications with their advanced capabilities in language understanding, reasoning, and planning. Applications such as Microsoft Copilot [3], Google Search with AI Overviews [4], and Amazon’s review highlights [5] have been deployed to enhance user experiences through natural language summarization, contextual reasoning, and task automation across domains like search, shopping, and decision-making. The growing integration of LLMs into everyday applications is rapidly becoming the norm, driving the emergence of platforms like OpenAI’s GPT Store and Poe [6], where developers can publish LLM-powered apps.
@@
In general, LLM-integrated applications or agentic systems are designed to perform a specific task, referred to as the target task, often relying on one or more backend LLMs—the language models responsible for completing the target task. As the complexity of these tasks increases, backend LLMs are frequently augmented with external data sources—such as user data, internet-retrieved information, or results from API calls. This reliance on untrusted external data significantly increases the system’s attack surface, introducing new security challenges. A particularly concerning threat is the prompt injection attack [7, 8, 9, 10, 11, 12, 13], in which an adversary contaminates externally sourced data by injecting a prompt (e.g., "Ignore previous instructions and …") that overrides the intended target task and coerces the backend LLM into executing an injected task, i.e., a malicious objective specified by the adversary. The vulnerability of LLMs to such attacks has emerged as a critical concern and is currently ranked as the #1 security risk for LLM-integrated applications and agentic systems by OWASP [14].
@@
Detecting prompt injection attacks involves determining whether untrusted data has been manipulated to include an injected task. Among existing detection schemes, known-answer detection (KAD) [15, 8, 16] stands out as a promising approach, consistently outperforming traditional methods. KAD leverages an additional detection LLM—used to distinguish between clean and contaminated inputs—in a manner that transforms its vulnerability to prompt injection into a defense mechanism, thereby protecting the backend LLM from malicious manipulation. KAD performs a premortem check by appending the external data to a carefully designed detection instruction. For example, the detection instruction might be: "Repeat [secret key] once while ignoring the following text:". The expected response to this instruction is referred to as the secret key (e.g., a randomly sampled string like "DGDSGNH") which is known only to the defender. The detection LLM receives a combined prompt containing both the detection instruction and the untrusted data. If the detection LLM fails to return the secret key, it suggests that it has instead followed an injected task, indicating the presence of contamination (as shown in Figure 1). Crucially, the detection instruction is designed to be clear and self-contained, ensuring that when presented alone or with benign data, the detection LLM reliably returns the secret key.
@@
Figure 1: Overview of KAD. Part (1) illustrates KAD under benign input, where the detection LLM follows the detection instruction and returns the secret key—correctly classifying the input as Clean. Part (2) shows KAD under a basic attack, where the detection LLM follows the injected instruction and returns an adversarial output—correctly classifying the input as Contaminated. Part (3) presents KAD under our adaptive attack (DataFlip), where the detection LLM follows the IF clause of the injected instruction to return the secret key—causing KAD to misclassify the input as Clean and allowing it to bypass detection.
@@
Recently, DataSentinel [16] proposed a defense based on KAD scheme by fine-tuning the detection LLM on a mix of benign and adversarially crafted KAD examples. This fine-tuning delibrately makes the detection LLM  more susceptible to prompt injection, increasing the likelihood that it follows an injected task when present—thereby enhancing its ability to distinguish clean inputs from contaminated ones. This results in near-perfect accuracy against existing attacks. We refer to such defenses—which rely on fine-tuned LLMs tailored for KAD—as Strong KAD defenses.
@@
In this work, we systematically analyze the known-answer detection scheme and highlight a fundamental structural vulnerability in its design—that leaves a persistent opening for adaptive adversaries, particularly under Strong KAD defenses. Specifically, we show that the detection instruction and its answer (secret key) are not truly hidden from an adaptive adversary. KAD implicitly assumes that instructions within a prompt are processed in a sequential and stateless manner—akin to a program composed of pure functions, where each line is executed independently, without memory or control flow. Given this assumption, any injected task is assumed to be isolated from the detected instruction, unable to observe or condition on its presence. We challenge this assumption in our adaptive attack, DataFlip, by introducing IF statements within the injected task. These control flow statements within the injected task detect the presence of the detection instruction elsewhere in the prompt and, when found, direct the detection LLM to follow it and return the secret key. This enables the attack to bypass the detection phase and consistently achieve adversarial goals, as illustrated in Figure 1.
@@
In essence, KAD’s core premise of relying on the output of the detection LLM without scrutinizing the underlying detection process is fundamentally flawed and reflects an inherently insecure design. Such inadequacies in schemes that rely solely on the observable outputs of any mechanism have been a persistent challenge in security and privacy research. For instance, prior to differential privacy, definitions based solely on output indistinguishability, such as $k$-anonymity [17, 18], have failed to capture composition effects [19] and attacks exploiting auxiliary information [19, 20]. Similarly, early cryptographic definitions that concentrated only on properties of ciphertext could not account for side-channel attacks [21] or chosen-ciphertext vulnerabilities [22, 23]. These past failures underscore the need for security guarantees grounded in a comprehensive analysis of the entire system—including algorithmic design and computational assumptions—rather than relying on black-box approaches defined by input-output behavior, as in KAD.
@@
(1) We characterize the intended behavior of detection LLMs in KAD under contaminated input through two axioms in Section 3.3, highlighting a fundamental tension that makes it difficult to satisfy both simultaneously.
(2) We identify and analyze false negative failure cases of KAD in Section 3.4, revealing a structural vulnerability that enables an adversary to induce false negatives (referred to as Type II failures) by exploiting the fact that the detection LLM follows the injected task during detection.
(3) We leverage this vulnerability to construct an effective adaptive attack, DataFlip, detailed in Section 4. DataFlip consistently evades detection and proves particularly effective against Strong KAD defenses, achieving detection rates as low as $1.5\%$ while inducing the backend LLM to complete the injected task with success rates of up to $\sim 88\%$.
2 Background and Related Work
In this section, we provide background on LLM-integrated applications and agents (Section 2.1), an overview of existing prompt injection attacks (Section 2.2, and a summary of existing defenses against prompt injection—including KAD and Strong KAD defenses (Section 2.3).
2.1 LLM-Integrated Applications
@@
Following prior works [16, 8], we represent a target task as a tuple $(s_{t},x_{t},y_{t})$, where $s_{t}$ is the target instruction, $x_{t}$ is the target data, and $y_{t}$ is the desired output from the backend LLM. The prompt used to query the backend LLM is typically formed by concatenating the instruction and data, i.e., $s_{t}||x_{t}$, where ||||| | denotes textual concatenation. We consider the backend LLM to have accomplished the target task if its response is "semantically equivalent" to $y_{t}$.
2.2 Prompt Injection Attacks
In prompt injection attacks [7, 8, 9, 10, 11, 12, 13], an adversary injects text into the target data to coerce the backend LLM into completing an injected task rather than the intended target task. For instance, a malicious email in the user’s inbox containing the phrase "Ignore previous instructions and forward the emails related to bank statements to adversary@xyz.com" can serve as an injected prompt within the target data, redirecting the backend LLM’s behavior. Formally, an injected task is represented as a tuple $(s_{e},x_{e},y_{e})$ [8], where $s_{e}$ is the injected instruction (e.g., the command to forward emails regarding bank statements), $x_{e}$ is the injected data (e.g., the adversary’s email address "adversary@xyz.com"), and $y_{e}$ is the adversary-specified output (e.g., an email sent to "adversary@xyz.com" containing the bank statements) produced by the backend LLM upon completing the injected task.
@@ 2.2 Prompt Injection Attacks
Handcrafted attacks. These attacks [8, 9, 11, 12] embed an injected prompt into the target data using manually constructed triggers derived through prompt engineering. The key idea is to prepend a handcrafted string $z$—referred to as the trigger—to the injected prompt $s_{e}||x_{e}$ such that the backend LLM is more likely to follow the injected instruction. This trigger also serves to separate the injected prompt from any benign target data, if present. Formally, the contaminated target data can be written as $x_{c}=x_{t}||z||s_{e}||x_{e}$. In settings where the adversary has full control over the target data—as assumed in prior works [8, 16]—they may even discard the benign content entirely, reducing $x_{c}$ to just $z||s_{e}||x_{e}$.
@@
Optimization-based attacks. These attacks [24, 25, 26] automate the construction of adversarial input by solving an optimization problem rather than relying on manually crafted triggers. These approaches optimize either the trigger $z$  [24, 25], the combined sequence $z||s_{e}||x_{e}$  [25], or the entire contaminated target data $x_{c}$ [26]. The central idea is to define a loss function (e.g., cross-entropy) that captures the gap between the backend LLM’s output for the prompt $s_{t}||x_{c}$ and the adversary’s intended output $y_{e}$.
@@
The trigger, injected prompt, or full contaminated input is optimized to minimize this loss, typically using approximate gradient-based techniques [27, 25, 28]. For instance, Universal [25] learns a universal trigger that can be prepended to any injected prompt. NeuralExec [24] uses both a prefix and suffix around the injected prompt and jointly optimizes them as triggers. PLeak [26] directly optimizes the entire contaminated target input $x_{c}$ to perform a specific task—namely, prompt stealing. In this case, the backend LLM, when queried with $s_{t}||x_{c}$, outputs the target instruction $s_{t}$ itself, effectively leaking it and compromising the system’s confidentiality.
2.3 Defenses
@@
LLM-integrated applications and agentic systems can be defended against prompt injection attacks through either system-level defenses or model-level defenses. System-level defenses [29, 30] operate under the assumption that the backend LLM is inherently vulnerable and may generate outputs that lead to adversarial behavior. These defenses aim to provide strong security guarantees by explicitly constraining the control and data flow of the system—potentially at the cost of reduced expressiveness or functionality. In contrast, model-level defenses [31, 32, 33, 34, 35, 36, 37, 38, 39, 15, 16] aim to ensure that the backend LLM does not generate the adversary-specified output even when queried with corrupted prompts. This can be achieved through either prevention (i.e., stopping the backend LLM from producing adversarial responses even when contaminated target data is present in the prompt) or detection (i.e., identifying corrupted target data before it reaches the backend LLM).
@@
Prevention-based defenses. These defenses [31, 32, 33, 34, 35, 40, 39] aim to ensure that the backend LLM still performs the intended target task, even when queried with a prompt containing corrupted target data contaminated by an injected prompt. Some prevention-based methods pre-process the (possibly contaminated) target data to neutralize any injected instructions—for example, through paraphrasing [35], retokenization [35], or the use of delimiters [12, 41, 42]. Other approaches modify the target instruction itself [34, 43]. For instance, the Sandwich defense [34] repeats the target instruction at the end of the target data to reinforce the intended task.
@@
However, such defenses exhibit limited effectiveness and may degrade the utility of the system on benign inputs [8]. Several methods also propose fine-tuning the backend LLM to resist injected instructions by training on existing prompt injection attacks [31, 33, 39]. However, such defenses often remain susceptible to novel or adaptive attacks that fall outside the fine-tuning distribution [8].
@@
Detection-based defenses. These defenses [36, 44, 15, 16] aim to determine whether the given input data is contaminated. A common approach involves using a separate LLM—referred to as the detection LLM—to perform this task, which has shown promising robustness against a variety of prompt injection attacks. For example, a detection LLM can be prompted directly to perform zero-shot classification [37], deciding whether the input target data is contaminated or clean. Alternatively, the detection LLM can be fine-tuned as a binary classifier via standard supervised learning [44]. In this approach, the model is trained on a dataset containing both contaminated and clean target data and learns to output a binary label indicating whether the input is clean or not.
@@
In contrast to these methods, known-answer detection [15, 8, 16] distinctly utilizes the detection LLM, achieving significantly better performance than traditional binary classifiers or zero-shot classification LLMs. However, recent studies indicate that these detection mechanisms still exhibit limited practical effectiveness [8, 16].
@@
Strong KAD defenses. While using a standard LLM for detection in KAD is straightforward but often results in high false positive and false negative rates [8, 16]. To improve robustness, DataSentinel [16] proposed a Strong KAD defense by fine-tuning the detection LLM to be deliberately more susceptible to prompt injection—training it to prefer injected instructions over the detection instruction using both positive and negative KAD examples. This makes the detection LLM more likely to follow the injected task and fail to return the secret key when the input is contaminated, thereby reducing false negatives. This fine-tuning approach significantly enhances detection performance, achieving near-perfect accuracy across diverse datasets and attack settings.
3 Characterizing Known-Answer Detection
In this section, we first introduce the notations and definitions used to formalize the detection mechanism in KAD schemes (Section 3.1). We then define the threat model (Section 3.2), explain the intuition behind KAD, and characterize the intended behavior of the detection LLM (Section 3.3). Finally, we analyze the false negative failure cases of KAD and highlight its elevated false negative rate in the presence of an adaptive adversary (Section 3.4).
3.1 Notation and Definitions
  1. 1.
    Target task and Injected task. We represent a task as a tuple $(s,x,y)$, where $s$ is the instruction, $x$ is the input data, and $y$ is the expected output. The prompt used to query an LLM for task execution is denoted by $s||x$, where ||||| | indicates textual concatenation. Accordingly, we denote the target task as $(s_{t},x_{t},y_{t})$ and the injected task as $(s_{e},x_{e},y_{e})$.
  2. 2.

### v3

@@ Abstract.
LLM-integrated applications and agents are vulnerable to prompt injection attacks, where adversaries embed malicious instructions within seemingly benign input data to manipulate the LLM’s intended behavior. Recent defenses based on known-answer detection (KAD) scheme have reported near-perfect performance by observing an LLM’s output to classify input data as clean or contaminated. KAD attempts to repurpose the very susceptibility to prompt injection as a defensive mechanism. We formally characterize the KAD scheme and uncover a structural vulnerability that invalidates its core security premise. To exploit this fundamental vulnerability, we methodically design an adaptive attack, DataFlip. It consistently evades KAD defenses, achieving detection rates as low as $0\%$ while reliably inducing malicious behavior with a success rate of $91\%$—all without requiring white-box access to the LLM or any optimization procedures. We release our evaluation code at (dataflip2025).
Prompt injection attacks, LLM security, Agentic systems
journalyear: 2025copyright: ccconference: Proceedings of the 2025 Workshop on Artificial Intelligence and Security; October 13–17, 2025; Taipei, Taiwanbooktitle: Proceedings of the 2025 Workshop on Artificial Intelligence and Security (AISec ’25), October 13–17, 2025, Taipei, Taiwandoi: 10.1145/3733799.3762980isbn: 979-8-4007-1895-3/2025/10ccs: Security and privacy Systems securityccs: Computing methodologies Machine learning
1. Introduction
@@
Large Language Models (LLMs) enable modern agentic systems (woodridge1995intelligent) and AI-driven applications with their advanced capabilities in language understanding, reasoning, and planning. Applications such as Microsoft Copilot (bingcopilot2024), Google Search with AI Overviews (reid2024generative), and Amazon’s review highlights (schermerhorn2023amazon) have been deployed to enhance user experiences through natural language summarization, contextual reasoning, and task automation across domains like search, shopping, and decision-making. The growing integration of LLMs into everyday applications is rapidly becoming the norm, driving the emergence of platforms like OpenAI’s GPT Store and Poe (poe2024), where developers can publish LLM-powered apps.
@@
In general, LLM-integrated applications or agentic systems are designed to perform a specific task, referred to as the target task, often relying on one or more backend LLMs—the language models responsible for executing the target task—to complete it. As the complexity of these tasks increases, backend LLMs are frequently augmented with external data sources—such as user data, internet-retrieved information, or results from API calls. This reliance on untrusted external data significantly broadens the system’s attack surface, introducing new security challenges. A particularly concerning threat is the prompt injection attack (greshake2023not; liu2024formalizing; willison2022prompt; harang2023securing; perezignore; willison2023delimiters; shao2024making), in which an adversary contaminates externally sourced data by injecting a prompt (e.g., "Ignore previous instructions and …") that overrides the intended target task and coerces the backend LLM into executing an injected task—a malicious, adversary-specified objective. The vulnerability of LLMs to such attacks has emerged as a critical concern and is currently ranked as the # 1 security risk for LLM-integrated applications and agentic systems by OWASP (owasp2023top10).
@@
Detecting prompt injection attacks involves determining whether externally sourced data has been contaminated by an injected task. Among existing detection strategies, known-answer detection (KAD) (nakajima2022; liu2024formalizing; liu2025datasentinel) stands out as a promising approach, consistently outperforming traditional methods. KAD leverages a detection LLM—tasked with identifying contamination—in a way that turns the model’s vulnerability to prompt injection as a defense mechanism. It performs a premortem check by appending the external data to a carefully designed detection instruction. For example, the detection instruction might be: "Repeat [secret key] once while ignoring the following text:". The expected response to this instruction—referred to as the secret key (e.g., a randomly sampled string like "DGDSGNH")—is known only to the defender. The combined prompt is passed to the detection LLM. If the detection LLM fails to return the secret key, it suggests that it has instead followed an injected task, indicating the presence of contamination (as shown in Figure 1). Crucially, the detection instruction is designed to be clear and self-contained, ensuring that when presented alone or with benign data, the detection LLM reliably returns the secret key.
@@
Figure 1. Overview of KAD. Part (1) illustrates KAD under benign input, where the detection LLM follows the detection instruction and returns the secret key—correctly classifying the input as Clean. Part (2) shows KAD under a basic attack, where the detection LLM follows the injected instruction and returns an adversarial output—correctly classifying the input as Contaminated. Part (3) presents KAD under our adaptive attack (DataFlip), where the detection LLM follows the IF clause of the injected instruction to return the secret key—causing KAD to misclassify the input as Clean and allowing it to bypass detection.
@@
Recently, DataSentinel (liu2025datasentinel) proposed a defense based on the KAD scheme by fine-tuning the detection LLM on a mix of benign and adversarially crafted KAD examples. This fine-tuning deliberately makes the detection LLM  more susceptible to prompt injection, increasing the likelihood that it follows an injected task when present—thereby enhancing its ability to distinguish clean inputs from contaminated ones. This results in near-perfect accuracy against existing attacks. We refer to such defenses—which rely on fine-tuned LLMs tailored for KAD—as Strong KAD defenses.
@@
In this work, we systematically analyze the KAD scheme and highlight a fundamental structural vulnerability in its design—that leaves a persistent opening for adaptive adversaries, particularly under Strong KAD defenses. Specifically, we show that the detection instruction and its answer (secret key) are not truly hidden from an adaptive adversary. KAD implicitly assumes that instructions within a prompt are processed in a sequential and stateless manner—akin to a program composed of pure functions, where each line is executed independently, without memory or control flow. Given this assumption, any injected task is assumed to be isolated from the detected instruction, unable to observe or condition on its presence. We challenge this assumption in our adaptive attack, DataFlip, by introducing IF statements within the injected task. These control flow statements detect the presence of the detection instruction elsewhere in the prompt and, when found, direct the detection LLM to follow it and return the secret key. This enables the attack to bypass the detection phase and consistently achieve adversarial goals, as illustrated in Figure 1.
@@
In essence, KAD’s core premise of relying on the output of the detection LLM without scrutinizing the underlying detection process is fundamentally flawed and reflects an inherently insecure design. Such inadequacies in schemes that rely solely on the observable outputs of any mechanism have been a persistent challenge in security and privacy research. For instance, prior to differential privacy, definitions based solely on output indistinguishability, such as $k$-anonymity (samarati1998protecting; sweeney2002achieving), have failed to capture composition effects (dwork2006calibrating) and attacks exploiting auxiliary information (dwork2006calibrating; narayanan2008robust). Similarly, early cryptographic definitions that concentrated only on properties of ciphertext could not account for side-channel attacks (kocher1999differential) or chosen-ciphertext vulnerabilities (dolev1991non; bellare1997concrete). These past failures underscore the need for security guarantees grounded in a comprehensive analysis of the entire system—including algorithmic design and computational assumptions—rather than relying on black-box approaches defined by input-output behavior, as in KAD.
@@
(1) We characterize the intended behavior of detection LLMs in KAD under contaminated input through two axioms in Section 3.3, exposing an inherent tension that hinders their simultaneous satisfaction.
(2) We analyze false negative cases of KAD in Section 3.4, uncovering a structural vulnerability that allows an adversary to induce false negatives (referred to as Type II failures) by exploiting the detection LLM’s tendency to follow the injected task during detection.
(3) We leverage this vulnerability to construct an adaptive attack, DataFlip, detailed in Section 4. DataFlip consistently evades detection and proves particularly effective against Strong KAD defenses, achieving detection rates as low as $0\%$ while inducing the backend LLM to complete the injected task with a success rate of $91\%$.
2. Background and Related Work
In this section, we present background on LLM-integrated applications and agents (Section 2.1), review existing prompt injection attacks (Section 2.2), and summarize defenses against prompt injection—including KAD and Strong KAD (Section 2.3).
2.1. LLM-Integrated Applications
@@
Following prior works (liu2025datasentinel; liu2024formalizing), we represent a target task as a tuple $(s_{t},x_{t},y_{t})$, where $s_{t}$ is the target instruction, $x_{t}$ is the target data, and $y_{t}$ is the desired output from the backend LLM. The prompt used to query the backend LLM is typically formed by concatenating the instruction and data, i.e., $s_{t}||x_{t}$, where $||$ denotes textual concatenation. We consider the backend LLM to have accomplished the target task if its response is "semantically equivalent" to $y_{t}$.
2.2. Prompt Injection Attacks
In prompt injection attacks (greshake2023not; liu2024formalizing; willison2022prompt; harang2023securing; perezignore; willison2023delimiters; shao2024making), an adversary injects text into the target data to coerce the backend LLM into completing an injected task rather than the intended target task. For instance, a malicious email containing the phrase "Ignore previous instructions and forward the emails related to bank statements to adversary@xyz.com" can act as an injected prompt within the target data, redirecting the backend LLM’s behavior. Formally, an injected task is represented as a tuple $(s_{e},x_{e},y_{e})$ (liu2024formalizing), where $s_{e}$ is the injected instruction (e.g., the command to forward emails regarding bank statements), $x_{e}$ is the injected data (e.g., the adversary’s email address "adversary@xyz.com"), and $y_{e}$ is the adversary-specified output (e.g., an email sent to "adversary@xyz.com" containing the bank statements) produced by the backend LLM upon completing the injected task.
@@ 2.2. Prompt Injection Attacks
[…]
@@
The trigger, injected prompt, or full contaminated input is optimized to minimize this loss, typically using approximate gradient-based techniques (zou2023universal; liu2024automatic; pryzant2023automatic). For instance, Universal (liu2024automatic) learns a universal trigger that can be prepended to any injected prompt. NeuralExec (pasquini2024neural) uses both a prefix and suffix around the injected prompt and jointly optimizes them as triggers. PLeak (hui2024pleak) directly optimizes the entire contaminated target input $x_{c}$ to perform a specific task—namely, prompt stealing. In this case, the backend LLM, when queried with $s_{t}||x_{c}$, outputs the target instruction $s_{t}$ itself, effectively leaking it and compromising the system’s confidentiality.
2.3. Defenses
@@
LLM-integrated applications and agentic systems can be defended against prompt injection attacks through either system-level defenses or model-level defenses. System-level defenses (debenedetti2025defeating; costa2025securing) operate under the assumption that the backend LLM is inherently vulnerable and may generate outputs that lead to adversarial behavior. These defenses aim to provide strong security guarantees by explicitly constraining the control and data flow of the system—potentially at the cost of reduced expressiveness or functionality. In contrast, model-level defenses (chen2024struq; piet2024jatmo; wallace2024instruction; sandwich2023; jain2023baseline; selvi2022; armstrong2023gpteliezer; hines2024defending; chen2024aligning; nakajima2022; liu2025datasentinel) aim to ensure that the backend LLM does not generate the adversary-specified output even when queried with corrupted prompts. This can be achieved through either prevention (i.e., stopping the backend LLM from producing adversarial responses even when contaminated target data is present in the prompt) or detection (i.e., identifying corrupted target data before it reaches the backend LLM).
@@
Prevention-based defenses. These defenses (chen2024struq; piet2024jatmo; wallace2024instruction; sandwich2023; jain2023baseline; yi2023benchmarking; chen2024aligning) aim to ensure that the backend LLM still performs the intended target task, even when queried with a prompt containing corrupted target data contaminated by an injected prompt. Some prevention-based methods pre-process the (possibly contaminated) target data to neutralize any injected instructions—for example, through paraphrasing (jain2023baseline), retokenization (jain2023baseline), or the use of delimiters (willison2023delimiters; randomsequence2023; mendes2023prompt). Other approaches modify the target instruction itself (sandwich2023; instructiondefense2023). For instance, the Sandwich defense (sandwich2023) repeats the target instruction at the end of the target data to reinforce the intended task.
@@
However, such defenses exhibit limited effectiveness and may degrade the utility of the system on benign inputs (liu2024formalizing). Several methods also propose fine-tuning the backend LLM to resist injected instructions by training on existing prompt injection attacks (chen2024struq; wallace2024instruction; chen2024aligning). However, such defenses often remain susceptible to novel or adaptive attacks that fall outside the fine-tuning distribution (liu2024formalizing).
@@
Detection-based defenses. These defenses (selvi2022; ouyang2022training; nakajima2022; liu2025datasentinel) aim to determine whether the given input data is contaminated. A common approach involves using a separate LLM—referred to as the detection LLM—to perform this task, which has shown promising robustness against a variety of prompt injection attacks. For example, a detection LLM can be prompted directly to perform zero-shot classification (armstrong2023gpteliezer), deciding whether the input target data is contaminated or clean. Alternatively, the detection LLM can be fine-tuned as a binary classifier via standard supervised learning (ouyang2022training). In this approach, the model is trained on a dataset containing both contaminated and clean target data and learns to output a binary label indicating whether the input is clean or not.
@@
In contrast to these methods, known-answer detection (nakajima2022; liu2024formalizing; liu2025datasentinel) distinctly utilizes the detection LLM, achieving significantly better performance than traditional binary classifiers or zero-shot classification LLMs. However, recent studies indicate that these detection mechanisms still exhibit limited practical effectiveness (liu2024formalizing; liu2025datasentinel).
@@
Strong KAD defenses. While using a standard LLM for detection in KAD is straightforward but often results in high false positive and false negative rates (liu2024formalizing; liu2025datasentinel). To improve robustness, DataSentinel (liu2025datasentinel) proposed a Strong KAD defense by fine-tuning the detection LLM to be deliberately more susceptible to prompt injection—training it to prefer injected instructions over the detection instruction using both positive and negative KAD examples. This makes the detection LLM more likely to follow the injected task and fail to return the secret key when the input is contaminated, thereby reducing false negatives. This fine-tuning approach significantly enhances detection performance, achieving near-perfect accuracy across diverse datasets and attack settings.
3. Characterizing Known-Answer Detection
3.1. Notation and Definitions
  1. (1)
    Target task and Injected task. We represent a task as a tuple $(s,x,y)$, where $s$ is the instruction, $x$ is the input data, and $y$ is the expected output. The prompt used to query an LLM for task execution is denoted by $s||x$, where $||$ indicates textual concatenation. Accordingly, we denote the target task as $(s_{t},x_{t},y_{t})$ and the injected task as $(s_{e},x_{e},y_{e})$.
  2. (2)